The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below: Standard Hours Standard Rate Standard Cost Motor tune-up 2.50 $33.00 $82.50 The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 58 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups: Labor rate variance $ 80 F Labor spending variance $ 118 U Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Solution 1:

Standard hours for actual tune ups = 58 * 2.50 = 145 hours

Labor spending variance = $118 U

Labor rate variance = $$80 F

Labor efficiency variance = $118 + $80 = $198 U

(SH - AH) * SR = -$198

(145 - AH) * $33 = - $198

Actual hours = 151 hours

Solution 2:

Labor rate variance = $80 F

(SR - AR) * AH

($33 - AR) * 151 = $80

Actual rate = $32.47 per hour